OpenStudy (anonymous):

What is the value of the expression when a = –6 and b = 5? (a+4b)^2+2a^2 A. –1061 B. 124 C. 268 D. –311

OpenStudy (vera_ewing):

so what do you think would be the first step?

OpenStudy (anonymous):

i already did the equation and i got 340 but thats not an answer choice.

OpenStudy (vera_ewing):

plug in -6 for a and 5 for b

OpenStudy (anonymous):

i know, i did that.

OpenStudy (vera_ewing):

(-6+4(5))^2+(2(-6))^2

OpenStudy (the_fizicx99):

Just plug in, (a+4b)^2+2a^2 (-6+4(5))^2+2(-6)^2 (-6 + 20)^2 + -12^2 -6^2 + 20^2 + 144 36 + 400 + 144 = 580

OpenStudy (the_fizicx99):

Does that make sense @queenleviosa

OpenStudy (vera_ewing):

(-6+20)^2+(-12)^2

OpenStudy (vera_ewing):

(14)^2+(-12)^2

OpenStudy (anonymous):

yes i know how to do the problem but the answers i get are not in the answer choices.

OpenStudy (vera_ewing):

196+144=?

OpenStudy (the_fizicx99):

I see the mistake you did you added -6 + 20 = 14 14^2 = 196 The thing is that you have to distribute the ^2 to both numbers inside the parenthesis.

OpenStudy (vera_ewing):

oh wow so i got 340 and it is not one of the choices! hmm let me try again hold on

OpenStudy (the_fizicx99):

Is 268 an option? I made a typo.

OpenStudy (vera_ewing):

ok so the answer is 268 :)

OpenStudy (vera_ewing):

oh like @tHe_FiZiCx99 said lol

OpenStudy (anonymous):

The answer is C. 268.

OpenStudy (anonymous):

how did u get that answer?

OpenStudy (vera_ewing):

you must distribute the ^2 to each term

OpenStudy (anonymous):

oh ok i got it now, thanks.

OpenStudy (vera_ewing):

yep np :)

OpenStudy (vera_ewing):

medal?